90 kN/m 165 KN 120 KN.m 3m 1m 2m А в 1mA 2m с D E F

Determine the moment and reaction at fix support. use any method.
Answer #1

P9-) given figure can be modified into Loading given below. F 180KN/m 165 KN 120kN.m 90 kN/m B D E Eim- 2m 3m VB 180kN/m Noweg-2 S LØPE at Deflection at B due to 165 Kn How shifting 105 at B we have moment of 165KNM and Load of 165KN. therefore 165x129-3 deflection at B due to VB 1 PL3 ЗЕТ V. BX 83 = 512VB (1) ЗЕТ 3 EI t we deflection at B due to upward traingolar LoadengВ due to Leg-deflection at Couple of 120KNM + 1 Cya аст t (а ох6 Х Таох 62 ДЕТ E Збоо (1) - Т.as per compatiblity Pa-5, AB = deflection at B=0 hence О= Дs - + 38160 t 5280 +то 6 EI EI EI 136 ОО - 512 MB — 1998 - 5 ЗЕТ ЄPg - 6 and another compatiblity We have 1 UU [VF + V₂ + VE = 165+ (135 kn *] KN . another compatiblity ean-e AE=0 + CE=0 165xlig-7 from ean (a) and (6) We have (VE = 66.65KN Putting value of VE in (b) (VB = 271.75KN) from ean-c Reaction at 5 = VE Sup